The amount of money that each person must contribute to buy a retirement gift for a coworker varies inversely as the number of people contributing. If 64 people must contribute ​$5 ​each, how much does each person have to contribute if there are 8 ​people.
Answer by Boreal(15235)

y=k/x; y=money, x=people
64=k/5
k=320 Use that in second part
y=320/8=$40 per person.